Circular No. 8503

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005az IN NGC 4960 R. Quimby, P. Mondol, P. Hoeflich, and J. C. Wheeler, University of

Texas; and C. Gerardy, Imperial College, report the discovery of an apparent supernova in unfiltered CCD images taken on Mar. 28.21 (at mag ∼ 17.4) and 31.45 UT (mag ∼ 16.9) using the 0.45-m ‘ROTSE-IIIb’ telescope at the McDonald Observatory. The new object is located at α = 13^{h}05^{m}46^{s}.97, δ = +27^{o}44′08′′.4 (equinox 2000.0), which is 8′′.0 west and 5′′.5 north of the center of NGC 4960. A co-addition of ROTSE-IIIb images taken between 2004 Dec. 15 and 2005 Jan. 14 (limiting mag ∼ 18.6) shows nothing at this position. The new source is not present on ROTSE-IIIb images taken on Mar. 20.23 (limiting mag ∼ 17.4).

SUPERNOVA 2005ba IN NGC 3746 S. Kildahl, Hokksund, Norway, reports the discovery by A. Danielsen

(Langhus, Norway), M. Steine (Blaker, Norway), and himself of an apparent supernova (mag ∼ 17.5) on unfiltered images taken with a ‘Celestron 14’ reflector at Veggli, Norway, on Apr. 1.92, 2.84, and 3.062 UT. The new object is located at α = 11^{h}37^{m}42^{s}.58, δ = +22^{o}00′38′′.9 (equinox 2000.0), which is 14′′.6 west and 4′′.0 north of the nucleus of NGC 3746. SN 2005ba was not detected in an earlier image taken by O. Trondal with the Tenagra II 0.81-m telescope on 2004 Dec. 12.38 (limiting mag ∼ 19.0).

SUPERNOVAE 2005al AND 2005au M. Salvo and B. Schmidt, Australian National University (ANU); and

J. Christiansen, University of New South Wales, write that a spectrum (range 407–570 nm) of SN 2005al (cf. IAUC 8488), taken with the ANU 2.3m telescope (+ Double-Beam Spectrograph) at Siding Spring Observatory on Feb. 28.59 UT, shows it to be a type-Ia supernova around maximum light.

M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smithsonian Center for Astrophysics; and T. Matheson, National Optical Astronomy Observatory, report that a spectrogram (range 350–740 nm) of SN 2005au (cf. IAUC 8496), obtained by P. Nutzman on Mar. 31.27 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), shows it to be a type-II supernova, probably within a few weeks past explosion. The spectrum consists of a blue continuum and P-Cyg lines of Hα and Hβ. Adopting the NED recession velocity of 5592 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is about 9000 km/s.

2005 April 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green